OpenStudy (anonymous):

The scores on a university examination are Normally distributed with a mean of 62 and a standard deviation of 11. If the bottom 5% of students will fail the course, what is the lowest mark that a student can have and still be awarded a passing grade? A) 62 B) 57 C) 44 D) 40 how do i get this answer?

Directrix (directrix):

C) 44 is what I cranked out. z = (x - 62)/11 5% is associate with a z-score of -1.64 -1.64 = (x-62)/11 x = 43.96